# Recent Advances in OOD Detection: Problems and Approaches

### v2

@@ Abstract.
Out-of-distribution (OOD) detection aims to detect test samples outside the training category space, which is an essential component in building reliable machine learning systems. Existing reviews on OOD detection primarily focus on method taxonomy, surveying the field by categorizing various approaches. However, many recent works concentrate on non-traditional OOD detection scenarios, such as test-time adaptation, multi-modal data sources and other novel contexts. In this survey, we uniquely review recent advances in OOD detection from the problem scenario perspective for the first time. According to whether the training process is completely controlled, we divide OOD detection methods into training-driven and training-agnostic. Besides, considering the rapid development of pre-trained models, large pre-trained model-based OOD detection is also regarded as an important category and discussed separately. Furthermore, we provide a discussion of the evaluation scenarios, a variety of applications, and several future research directions. We believe this survey with new taxonomy will benefit the proposal of new methods and the expansion of more practical scenarios. A curated list of related papers is provided in the Github repository: https://github.com/shuolucs/Awesome-Out-Of-Distribution-Detection
@@ 1. Introduction
Machine learning methods  (LeCun et al., 2015) have made significant progress under the closed-world assumption, where test data is drawn from the same distribution as the training set, known as in-distribution (ID). However, in the real world, models inevitably encounter test samples that do not belong to any training set category, commonly referred to as out-of-distribution (OOD) data. OOD detection  (Hendrycks and Gimpel, 2017) aims to identify and reject OOD samples rather than make overconfident predictions arbitrarily  (Nguyen et al., 2015) while maintaining accurate classification for ID data. Models with superior OOD detection capabilities are more reliable and have important applications in numerous security-critical scenarios. For instance, in medical diagnosis systems, a model that cannot detect OOD samples will misjudge unknown diseases and cause serious misdiagnosis  (Schlegl et al., 2017). Similarly, autonomous driving algorithms  (Geiger et al., 2012) should detect unknown scenarios and resort to human control to avoid accidents caused by arbitrary judgment.
@@
Notably, several previous efforts have been dedicated to surveying and summarizing OOD detection in recent years.  Yang et al. (2021) discuss OOD detection with several similar topics and categorize existing work into classification-based, density-based, distance-based, and reconstruction-based methods.  Cui and Wang (2022) conduct a survey on OOD detection from a methodological perspective but with an alternative classification criterion, including supervised, semi-supervised, and unsupervised methods. Additionally,  Lang et al. (2023) offer a review of OOD detection methods in natural language processing. However, previous works focus too much on the discussion from the perspective of methods and lack an in-depth exploration from the viewpoint of task scenarios. Establishing a clear taxonomy of task scenarios can enhance a comprehensive understanding of the field and assist practitioners in selecting the appropriate method. Moreover, given the recent introduction of new paradigms (e.g., test-time learning paradigm  (Yang et al., 2023a)) and methods based on large pre-trained models  (Esmaeilpour et al., 2022), there is an urgent need for a survey that incorporates the latest technologies.
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=7pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=14em,font=,, where level=2text width=13em,font=,, where level=3text width=11em,font=,, where level=4text width=10em,font=,, [ OOD Detection ,fill=hidden-yellow!70,ver [ Problem: Training-driven
@@
OOD detection(§ 3),fill=hidden-red!70, align=center, text width=12em [ Approaches with only
@@
ID Data (§ 3.1),fill=hidden-blue!70,align=center [ Reconstruction-based, fill=hidden-blue!70 [ MoodCat  (Yang et al., 2022a), RAE (Zhou, 2022), MOOD  (Li et al., 2023a), MOODv2  (Li et al., 2024a), PRE  (Osada et al., 2023), LMD  (Liu et al., 2023a) DiffGuard  (Gao et al., 2023a),
@@
DenoDiff (Graham et al., 2023) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ Probability-based, fill=hidden-blue!70 [ HVCM (Li et al., 2023b), DDR (Huang et al., 2022), LID (Kamkari et al., 2024) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ Logits-based, fill=hidden-blue!70 [ LogitNorm  (Wei et al., 2022a), UE-NL  (Huang and Qiao, 2023), DML  (Zhang and Xiang, 2023) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ OOD Synthesis, fill=hidden-blue!70 [ ConfiCali (Lee et al., 2018a), CODEs  (Tang et al., 2021), CMG  (Wang et al., 2022a), VOS  (Du et al., 2022a), NPOS  (Tao et al., 2023), SHIFT  (Kwon et al., 2023), ATOL  (Zheng et al., 2023),
@@
SSOD  (Pei, 2024), SEM  (Yang et al., 2023b) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ Prototype-based, fill=hidden-blue!70 [ SIREN  (Du et al., 2022b), CIDER  (Ming et al., 2023), PALM  (Lu et al., 2024), ReweightOOD  (Regmi et al., 2024), Step  (Zhou et al., 2021) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ A special case:
@@
Long-tail ID data, fill=hidden-blue!70 [ OLTR  (Liu et al., 2019), II-Mixup (Mehta et al., 2022), AREO  (Sapkota and Yu, 2023), IDCP (Jiang et al., 2023a), Open-Sampling  (Wei et al., 2022b), COOD  (Hogeweg et al., 2024) , leaf, text width=38em,fill=hidden-blue!70 ] ] ] [ Approaches with both ID
@@
and OOD Data (§ 3.2),fill=hidden-blue!70,align=center [ Boundary Regularization,fill=hidden-blue!70 [ OE  (Hendrycks et al., 2019), ELOC (Vyas et al., 2018), Why-ReLU (Hein et al., 2019), SSL-GOOD (Mohseni et al., 2020), EnergyOE  (Liu et al., 2020), MixOE  (Zhang et al., 2023a) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ Outlier Mining,fill=hidden-blue!70 [ BD-Resamp (Li and Vasconcelos, 2020), POEM  (Ming et al., 2022a), DAOL  (Wang et al., 2023a), DOE (Wang et al., 2023b), MixOE  (Zhang et al., 2023a), DivOE  (Zhu et al., 2023) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ Imbalanced ID,fill=hidden-blue!70 [ PASCAL  (Wang et al., 2022b), COCL  (Miao et al., 2024), BERL (Choi et al., 2023), EAT  (Wei et al., 2024) , leaf, text width=38em,fill=hidden-blue!70 ] ] ] ] [ Problem: Training-agnostic
@@
OOD detection (§ 4),fill=hidden-red!70,align=center,text width=12em [ Post-hoc Approaches (§ 4.1),fill=hidden-blue!70,align=center [ Output-based,fill=hidden-blue!70 [ MSP  (Hendrycks and Gimpel, 2017), MaxLogits  (Hendrycks et al., 2022), Energy  (Liu et al., 2020), GEN  (Liu et al., 2023b), ZODE  (Xue et al., 2024), LogicOOD  (Kirchheim et al., 2024) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ Distance-based,fill=hidden-blue!70 [ Mahalanobis  (Lee et al., 2018b), NNGuide  (Park et al., 2023a), KNN  (Sun et al., 2022), SSD  (Sehwag et al., 2021) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ Gradient-based,fill=hidden-blue!70 [ Grad  (Lee and AlRegib, 2020), GradNorm  (Huang et al., 2021), GradOrth  (Behpour et al., 2023), GAIA  (Chen et al., 2023), OPNP  (Chen et al., 2024a) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ Feature-based,fill=hidden-blue!70 [ ODIN  (Liang et al., 2018a), ReAct  (Sun et al., 2021), VRA  (Xu et al., 2023), SHE  (Zhang et al., 2022), ViM  (Wang et al., 2022c), Neco  (Ammar et al., 2024), ASH  (Djurisic et al., 2023), NAC  (Liu et al., 2024a)
@@
Optimal-FS (Zhao et al., 2024), BLOOD  (Jelenić et al., 2024), SCALE  (Xu et al., 2024), DDCS  (Yuan et al., 2024), LINe  (Yong Hyun Ahn, 2023) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ Density-based,fill=hidden-blue!70 [ GEM  (Morteza and Li, 2022), ConjNorm  (Peng et al., 2024) , leaf, text width=38em,fill=hidden-blue!70 ] ] ] [ Test-time Adaptive Approaches
@@
(§ 4.2) ,fill=hidden-blue!70,align=center [ Model-update needed,fill=hidden-blue!70 [ WOODS  (Katz-Samuels et al., 2022), AUTO  (Yang et al., 2023a), SODA  (Geng et al., 2023), ATTA  (Gao et al., 2023b), SAL  (Du et al., 2024) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ Model-update free,fill=hidden-blue!70 [ ETLT  (Fan et al., 2022), AdaOOD  (Zhang et al., 2023b), GOODAT  (Wang et al., 2024), RTL  (Fan et al., 2024) , leaf, text width=38em,fill=hidden-blue!70 ] ] ] ] [ Problem: Large pre-trained
@@
(§ 5),fill=hidden-red!70,align=center,text width=12em [ Zero-shot Approaches
@@
(§ 5.1),fill=hidden-blue!70,align=center [ VLM-based,fill=hidden-blue!70 [ CLIPScope  (Fu et al., 2024), RONIN  (Nguyen et al., 2024), ZOC  (Esmaeilpour et al., 2022), MCM  (Ming et al., 2022b), CLIPN  (Wang et al., 2023c), NegLabel  (Jiang et al., 2023b),
@@
NegPrompt-C (Nie et al., 2024), LAPT  (Zhang et al., 2024) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ LLM-based,fill=hidden-blue!70 [ WK-LLM (Dai et al., 2023), ODPC  (Huang et al., 2024), LLM-OOD (Liu et al., 2024b) , leaf, text width=38em,fill=hidden-blue!70 ] ] ] [ Few-shot Approaches
@@
(§ 5.2),fill=hidden-blue!70,align=center [ Fine-tuning based, fill=hidden-blue!70 [ CLIP-OS  (Sun et al., 2024), EOK-Prompt  (Zeng et al., 2024), GalLop  (Lafon et al., 2024), ID-like  (Bai et al., 2024), DSFG  (Dong et al., 2023), LoCoOp  (Miyai et al., 2023a),
@@
NegPrompt  (Li et al., 2024b), TagOOD  (Li et al., 2024c), CRoFT (Zhu et al., 2024) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ Meta-learning based, fill=hidden-blue!70 [ OOD-MAML  (Jeong and Kim, 2020), HyperMix  (Mehta et al., 2024) , leaf, text width=38em,fill=hidden-blue!70 ] ] ] [ Full-shot Approaches
@@
(§ 5.3),fill=hidden-blue!70,align=center [ NPOS  (Tao et al., 2023), PT-OOD (Miyai et al., 2023b), TOE  (Park et al., 2023b) , leaf, text width=48.5em,fill=hidden-blue!70,text width=50.6em ] ] ] ]
Figure 1. Taxonomy of OOD detection problem scenarios and solutions.
@@
In this survey, we for the first time review the recent advances in OOD detection with a problem-oriented taxonomy, as illustrated in Fig. 1. Based on whether the method needs to control the pre-training process, we categorize OOD detection algorithms into training-driven and training-agnostic methods. Considering the rapid development of large pre-trained models nowadays, we also regard large pre-trained model-based OOD detection as a separate section. In detail, training-driven methods achieve high detection capability by designing the optimization process of the training stage. They are further classified and discussed according to whether OOD data is used in training. Training-agnostic methods distinguish OOD data from ID ones based on a well-trained model, skipping the time-consuming and expensive pre-training process in practice. According to whether utilizing test samples to further improve OOD detection performance, we categorize them into post-hoc and test-time methods. Large pre-trained model-based OOD detection methods focus on models such as vision language models or large language models, which are pre-trained on vast datasets and excel in a wide array of tasks. We discuss them in terms of whether they have access to a few examples, including zero-shot, few-shot and full-shot scenarios.
@@
The remainder of this survey is organized as follows. We recap the related work of OOD detection in Sec. 2. Next, we summarize training-driven OOD detection approaches in Sec. 3, and introduce the training-agnostic OOD detection methods in Sec. 4. Then, in Sec. 5, we introduce large pre-trained model-based OOD detection. An overview of the evaluation metrics, experimental protocols, and applications is presented in Sec. 6. Following that, we discuss promising trends and open challenges in Sec. 7 to shed light on underexplored and potentially critical avenues. Finally, we conclude this survey in Sec. 8.

### v3

# Out-of-Distribution Detection: A Task-Oriented Survey of Recent Advances

@@ Abstract.
Out-of-distribution (OOD) detection aims to detect test samples outside the training category space, which is an essential component in building reliable machine learning systems. Existing reviews on OOD detection primarily focus on method taxonomy, surveying the field by categorizing various approaches. However, many recent works concentrate on non-traditional OOD detection scenarios, such as test-time adaptation, multi-modal data sources and other novel contexts. In this survey, we uniquely review recent advances in OOD detection from the task-oriented perspective for the first time. According to the user’s access to the model, that is, whether the OOD detection method is allowed to modify or retrain the model, we classify the methods as training-driven or training-agnostic. Besides, considering the rapid development of pre-trained models, large pre-trained model-based OOD detection is also regarded as an important category and discussed separately. Furthermore, we provide a discussion of the evaluation scenarios, a variety of applications, and several future research directions. We believe this survey with new taxonomy will benefit the proposal of new methods and the expansion of more practical scenarios. A curated list of related papers is provided in the Github repository: https://github.com/shuolucs/Awesome-Out-Of-Distribution-Detection
@@ 1. Introduction
Machine learning methods have made significant progress under the closed-world assumption, where test data is drawn from the same category as the training set, known as in-distribution (ID). However, in the real world, models inevitably encounter test samples that do not belong to any training set category, commonly referred to as out-of-distribution (OOD) data. OOD detection  (Hendrycks and Gimpel, 2017) aims to identify and reject OOD samples rather than make overconfident predictions arbitrarily  (Nguyen et al., 2015) while maintaining accurate classification for ID data. Models with superior OOD detection capabilities are more reliable and have important applications in numerous security-critical scenarios. For instance, in medical diagnosis systems, a model that cannot detect OOD samples will misjudge unknown diseases and cause serious misdiagnosis  (Schlegl et al., 2017). Similarly, autonomous driving algorithms  (Geiger et al., 2012) should detect unknown scenarios and resort to human control to avoid accidents caused by arbitrary judgment.
Table 1. Comparison of recent survey papers on OOD detection. “Type” indicates whether the survey is organized from a methodological perspective (“Method”) or from a broader task-oriented perspective (“Task”). “OOD-Only” indicates an exclusive focus on OOD detection; “LPMs” refers to coverage of methods based on large pretrained models; “Application” refers to in-depth analysis of real-world use cases; “Multi-Modal” indicates consideration of multiple data modalities; “Test-Time” indicates surveys that explicitly distinguish test-time adaptive methods from post-hoc approaches and provide dedicated discussion on test-time adaptation.
Study Year Type OOD-Only LPMs Application Multi-Modal Test-Time
Salehi et al. (Salehi et al., 2022) 2021 Method
Yang et al. (Yang et al., 2021) 2021 Method
Cui et al. (Cui and Wang, 2022) 2022 Method
Lang et al. (Lang et al., 2023) 2023 Method
Xu et al. (Xu and Ding, 2024) 2024 Method
Miyai et al. (Miyai et al., 2024) 2024 Method
Ours 2024 Task
@@
Notably, several previous efforts have been dedicated to surveying and summarizing OOD detection in recent years.  Salehi et al. (2022) offer a comprehensive review covering Anomaly Detection, Novelty Detection, Open-Set Recognition and OOD Detection, and analyzes the relationships and distinctions among these fields.  Yang et al. (2021) propose a unified framework to discuss OOD detection with several similar topics and categorize existing work into classification-based, density-based, distance-based and reconstruction-based methods. Concurrent with our work,  Miyai et al. (2024) discuss CLIP-based methods across related domains, while our survey specifically centers on OOD detection and provides a more comprehensive coverage of recent advances in this area.  Cui and Wang (2022) conduct a survey on OOD detection from a methodological perspective but with an alternative classification criterion, including supervised, semi-supervised, and unsupervised methods. Additionally,  Lang et al. (2023); Xu and Ding (2024) review various OOD detection methods in the context of natural language processing. However, previous works focus too much on the discussion from the perspective of methods and lack an in-depth exploration from the viewpoint of task scenarios. Establishing a clear taxonomy of task scenarios can enhance a comprehensive understanding of the field and assist practitioners in selecting the appropriate method. Moreover, given the recent introduction of new paradigms (e.g., test-time learning paradigm  (Yang et al., 2023a; Geng et al., 2023; Gao et al., 2023a; Du et al., 2024a; Fan et al., 2022; Wang et al., 2024a)) and methods based on large pre-trained models  (Esmaeilpour et al., 2022; Nguyen et al., 2024; Fu et al., 2024; Dai et al., 2023; Liu et al., 2024a), there is an urgent need for a comprehensive survey that incorporates the latest technologies.
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=7pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=14em,font=,, where level=2text width=13em,font=,, where level=3text width=11em,font=,, where level=4text width=10em,font=,, [ OOD Detection ,fill=hidden-red!70, draw=hidden-red!80!black,ver [ Problem: Training-driven
@@
OOD detection(§ 3),fill=hidden-blue!70, draw=hidden-blue!80!black, align=center, text width=12em [ Approaches with only
@@
ID Data (§ 3.1),fill=hidden-blue!50, draw=hidden-blue!80!black,align=center [ Reconstruction-based, fill=hidden-blue!30, draw=hidden-blue!80!black, [ MoodCat  (Yang et al., 2022), RAE (Zhou, 2022), MOOD  (Li et al., 2023a), MOODv2  (Li et al., 2024a), PRE  (Osada et al., 2023), LMD  (Liu et al., 2023a) DiffGuard  (Gao et al., 2023b),
@@
DenoDiff (Graham et al., 2023) , leaf, text width=38em,fill=hidden-blue!20, draw=hidden-blue!80!black, ] ] [ Probability-based, fill=hidden-blue!30, draw=hidden-blue!80!black, [ HVCM (Li et al., 2023b), DDR (Huang et al., 2022a), LID (Kamkari et al., 2024) , leaf, text width=38em,fill=hidden-blue!20, draw=hidden-blue!80!black, ] ] [ Logits-based, fill=hidden-blue!30, draw=hidden-blue!80!black, [ LogitNorm  (Wei et al., 2022a), UE-NL  (Huang and Qiao, 2023), DML  (Zhang and Xiang, 2023) , leaf, text width=38em,fill=hidden-blue!20, draw=hidden-blue!80!black, ] ] [ OOD Synthesis, fill=hidden-blue!30, draw=hidden-blue!80!black, [ ConfiCali (Lee et al., 2018a), CODEs  (Tang et al., 2021), CMG  (Wang et al., 2022a), VOS  (Du et al., 2022a), NPOS  (Tao et al., 2023), SHIFT  (Kwon et al., 2023), ATOL  (Zheng et al., 2023),
@@
SSOD  (Pei, 2024), SEM  (Yang et al., 2023b), Forte  (Ganguly et al., 2025), HamOS  (Li and Zhang, 2025), POP  (Gong et al., 2025) , leaf, text width=38em,fill=hidden-blue!20, draw=hidden-blue!80!black, ] ] [ Prototype-based, fill=hidden-blue!30, draw=hidden-blue!80!black, [ Step  (Zhou et al., 2021), SIREN  (Du et al., 2022b), CIDER  (Ming et al., 2023), PALM  (Lu et al., 2024), ReweightOOD  (Regmi et al., 2024), AROS  (Mirzaei and Mathis, 2025), PFS  (Wu et al., 2025) , leaf, text width=38em,fill=hidden-blue!20, draw=hidden-blue!80!black, ] ] [ A special case:
@@
Long-tail ID data, fill=hidden-blue!30, draw=hidden-blue!80!black, [ OLTR  (Liu et al., 2019), II-Mixup (Mehta et al., 2022), AREO  (Sapkota and Yu, 2023), IDCP (Jiang et al., 2023a), Open-Sampling  (Wei et al., 2022b), COOD  (Hogeweg et al., 2024) , leaf, text width=38em,fill=hidden-blue!20, draw=hidden-blue!80!black, ] ] ] [ Approaches with both ID
@@
and OOD Data (§ 3.2),fill=hidden-blue!50, draw=hidden-blue!80!black,align=center [ Boundary Regularization,fill=hidden-blue!30, draw=hidden-blue!80!black, [ OE  (Hendrycks et al., 2019), ELOC (Vyas et al., 2018), Why-ReLU (Hein et al., 2019), SSL-GOOD (Mohseni et al., 2020), EnergyOE  (Liu et al., 2020a), MixOE  (Zhang et al., 2023a) , leaf, text width=38em,fill=hidden-blue!20, draw=hidden-blue!80!black, ] ] [ Outlier Mining,fill=hidden-blue!30, draw=hidden-blue!80!black, [ BD-Resamp (Li and Vasconcelos, 2020), POEM  (Ming et al., 2022a), DAOL  (Wang et al., 2023a), DOE (Wang et al., 2023b), MixOE  (Zhang et al., 2023a), DivOE  (Zhu et al., 2023) , leaf, text width=38em,fill=hidden-blue!20, draw=hidden-blue!80!black, ] ] [ Imbalanced ID,fill=hidden-blue!30, draw=hidden-blue!80!black, [ PASCAL  (Wang et al., 2022b), COCL  (Miao et al., 2024a), BERL (Choi et al., 2023), EAT  (Wei et al., 2024) , leaf, text width=38em,fill=hidden-blue!20, draw=hidden-blue!80!black, ] ] ] ] [ Problem: Training-agnostic
@@
OOD detection (§ 4),fill=hidden-blue!70, draw=hidden-blue!80!black,align=center,text width=12em [ Post-hoc Approaches (§ 4.1),fill=hidden-blue!50, draw=hidden-blue!80!black,align=center [ Output-based,fill=hidden-blue!30, draw=hidden-blue!80!black, [ MSP  (Hendrycks and Gimpel, 2017), MaxLogits  (Hendrycks et al., 2022), Energy  (Liu et al., 2020a), GEN  (Liu et al., 2023b), ZODE  (Xue et al., 2024), LogicOOD  (Kirchheim et al., 2024) , leaf, text width=38em,fill=hidden-blue!20, draw=hidden-blue!80!black, ] ] [ Distance-based,fill=hidden-blue!30, draw=hidden-blue!80!black, [ Mahalanobis  (Lee et al., 2018b), NNGuide  (Park et al., 2023a), KNN  (Sun et al., 2022), SSD  (Sehwag et al., 2021),Mahalanobis++ (Mueller and Hein, 2025a) , leaf, text width=38em,fill=hidden-blue!20, draw=hidden-blue!80!black, ] ] [ Gradient-based,fill=hidden-blue!30, draw=hidden-blue!80!black, [ Grad  (Lee and AlRegib, 2020), GradNorm  (Huang et al., 2021), GradOrth  (Behpour et al., 2023), GAIA  (Chen et al., 2023), OPNP  (Chen et al., 2024a), PRO  (Chen et al., 2025a) , leaf, text width=38em,fill=hidden-blue!20, draw=hidden-blue!80!black, ] ] [ Feature-based,fill=hidden-blue!30, draw=hidden-blue!80!black, [ ODIN  (Liang et al., 2018a), ReAct  (Sun et al., 2021), VRA  (Xu et al., 2023), SHE  (Zhang et al., 2022), ViM  (Wang et al., 2022c), Neco  (Ammar et al., 2024), ASH  (Djurisic et al., 2023), NAC  (Liu et al., 2024b)
Optimal-FS (Zhao et al., 2024), BLOOD  (Jelenić et al., 2024), SCALE  (Xu et al., 2024a), DDCS  (Yuan et al., 2024), LINe  (Yong Hyun Ahn, 2023), KANs  (Canevaro et al., 2025),
@@
CADRef  (Ling et al., 2025), ITP  (Xu and Yang, 2025), NCI  (Liu and Qin, 2023) , leaf, text width=38em,fill=hidden-blue!20, draw=hidden-blue!80!black, ] ] [ Density-based,fill=hidden-blue!30, draw=hidden-blue!80!black, [ GEM  (Morteza and Li, 2022), ConjNorm  (Peng et al., 2024) , leaf, text width=38em,fill=hidden-blue!20, draw=hidden-blue!80!black, ] ] ] [ Test-time Adaptive Approaches
@@
(§ 4.2) ,fill=hidden-blue!50, draw=hidden-blue!80!black,align=center [ Model-update needed,fill=hidden-blue!30, draw=hidden-blue!80!black, [ WOODS  (Katz-Samuels et al., 2022), AUTO  (Yang et al., 2023a), SODA  (Geng et al., 2023), ATTA  (Gao et al., 2023a), SAL  (Du et al., 2024a) , leaf, text width=38em,fill=hidden-blue!20, draw=hidden-blue!80!black, ] ] [ Model-update free,fill=hidden-blue!30, draw=hidden-blue!80!black, [ ETLT  (Fan et al., 2022), AdaOOD  (Zhang et al., 2023b), GOODAT  (Wang et al., 2024a), RTL  (Fan et al., 2024), OODD  (Yang et al., 2025) , leaf, text width=38em,fill=hidden-blue!20, draw=hidden-blue!80!black, ] ] ] ] [ Problem: Large pre-trained
@@
(§ 5),fill=hidden-green!70,draw=hidden-green!80!black,align=center,text width=12em [ Zero-shot Approaches
@@
(§ 5.1),fill=hidden-green!50,draw=hidden-green!80!black,align=center [ VLM-based,fill=hidden-green!20,draw=hidden-green!80!black, [ CLIPScope  (Fu et al., 2024), RONIN  (Nguyen et al., 2024), ZOC  (Esmaeilpour et al., 2022), MCM  (Ming et al., 2022b), CLIPN  (Wang et al., 2023c), NegLabel  (Jiang et al., 2023b),
NegPrompt-C (Nie et al., 2024), LAPT  (Zhang et al., 2024a) , AdaNeg (Zhang and Zhang, 2024), CSP (Chen et al., 2024b) , SimLabel (Zou et al., 2025) ,  (Jung et al., 2024) ,
SeTAR (Li et al., 2024b) , OT-DETECTOR (Liu et al., 2025a) , leaf, text width=38em,fill=hidden-green!20,draw=hidden-green!80!black, ] ] [ LLM-based,fill=hidden-green!20,draw=hidden-green!80!black, [ WK-LLM (Dai et al., 2023), ODPC  (Huang et al., 2024), LLM-OOD (Liu et al., 2024a), CMA (Lee et al., 2025), ReGuide (Kim et al., 2025) , COOD (Liu et al., 2024c),
@@
EOE (Cao et al., 2024) , leaf, text width=38em,fill=hidden-green!20,draw=hidden-green!80!black, ] ] ] [ Few-shot Approaches
@@
(§ 5.2),fill=hidden-green!50,draw=hidden-green!80!black,align=center [ Fine-tuning based, fill=hidden-green!20,draw=hidden-green!80!black, [ CLIP-OS  (Sun et al., 2024), Local-Prompt  (Zeng et al., 2025), GalLop  (Lafon et al., 2024), ID-like  (Bai et al., 2024a), DSFG  (Dong et al., 2023), LoCoOp  (Miyai et al., 2023a),
@@
NegPrompt  (Li et al., 2024c), SUPREME (Wang et al., 2025a) , SCT (Yu et al., 2024a) , GaCoOp (Tong et al., 2025) , leaf, text width=38em,fill=hidden-green!20,draw=hidden-green!80!black, ] ] [ Meta-learning based, fill=hidden-green!20,draw=hidden-green!80!black,, [ OOD-MAML  (Jeong and Kim, 2020), HyperMix  (Mehta et al., 2024) , leaf, text width=38em,fill=hidden-green!20,draw=hidden-green!80!black, ] ] ] [ Full-shot Approaches
@@
(§ 5.3),fill=hidden-green!50,draw=hidden-green!80!black,align=center [ NPOS  (Tao et al., 2023), PT-OOD (Miyai et al., 2023b), TOE  (Park et al., 2023b) , leaf, text width=48.5em,fill=hidden-green!20,draw=hidden-green!80!black,,text width=50.6em ] ] ] ]
Figure 1. Taxonomy of OOD detection problem scenarios and solutions. The first two categories (in blue) correspond to shallow model-based methods, while the third category (in green) highlights approaches based on foundation models (large pre-trained models). Different colors are used to visually distinguish the traditional shallow-model-based solutions from the newly emerging foundation model-based methods, emphasizing the conceptual differences between them.
@@
In this survey, we for the first time review the recent advances in OOD detection with a task-oriented taxonomy, as illustrated in Fig. 1. Based on whether the method needs to control the pre-training process, we categorize OOD detection algorithms into training-driven and training-agnostic methods. Considering the rapid development of large pre-trained models nowadays and the evolving definition of OOD under such models, we also regard large pre-trained model-based OOD detection as a separate section. In detail, training-driven methods achieve high detection capability by designing the optimization process of the training stage. They are further classified and discussed according to whether OOD data is used in training. Training-agnostic methods distinguish OOD data from ID ones based on a well-trained model, skipping the time-consuming and expensive pre-training process in practice. According to whether utilizing test samples to further improve OOD detection performance, we categorize them into post-hoc and test-time methods. Large pre-trained model-based OOD detection methods focus on models such as vision language models or large language models, which are pre-trained on vast datasets and excel in a wide array of tasks. We discuss them in terms of whether they have access to a few examples, including zero-shot, few-shot and full-shot scenarios.
@@
The structure of this survey is as follows. We discuss the related topics of OOD detection in Sec. 2. Next, we summarize training-driven OOD detection approaches in Sec. 3, and introduce the training-agnostic OOD detection methods in Sec. 4. Then, in Sec. 5, we introduce large pre-trained model-based OOD detection. An overview of the evaluation metrics, experimental protocols, and applications is presented in Sec. 6. Following that, we discuss promising trends and open challenges in Sec. 7 to shed light on underexplored and potentially critical avenues.